# A Survey on Open Information Extraction from Rule-based Model to Large Language Model

### v6

@@ 6.2 Future Directions
Chronologically, IE systems increasingly employ diverse information sources and approaches, and now begin to converge on utilizing universal formats for various tasks (UIE). This shift is driven by advances in NLP techniques, especially pre-training models that enhance text extraction, and by increased computing power over the last two decades that enables more complex models. Recently, the exceptional language understanding and generalization capabilities of LLMs are promoting a shift towards UIE, broadening applications across different domains. More details are in Appendix C.
OpenIE datasets are growing but remain small relative to web information. Currently, these datasets are mostly confined to sources like Wiki, Newswire, NYT, and Freebase, with limited multilingual and multi-source corpus. Future expansions should include more languages and broader sources. Additionally, there is a need for synthesized datasets to improve both quality and quantity in OpenIE, as discussed in Appendix D. This could facilitate the creation of cross-domain datasets and integration of existing datasets and tasks.
Overly-specific relation output and the lack of a standard form for OpenIE output continue to challenge current models. Over-specificity in OpenIE arises from metrics focusing on token rather than semantic similarity, leading to verbose and incomplete outputs, while in ORC, sentences containing multiple relations introduce noise, complicating clustering and downstream tasks. Furthermore, the absence of standardized outputs hinders model comparison and canonicalization. Future research should focus on developing semantic-level evaluation metrics for OpenIE and establishing output standards tailored to downstream task requirements, alongside exploring text purification strategies for ORC to isolate distinct relations.
@@
Despite advances, most current LLM-based UIE systems focus on traditional IE tasks and often overlook OpenIE, a complex challenge within the IE spectrum. LLMs are inherently suited for OpenIE due to their extensive pre-trained knowledge, unlike smaller models that require extensive training to learn relational information. The primary challenge of LLMs lies not in extracting relational information but in accurately interpreting and following task-specific instructions.
@@
Our survey primarily concentrates on the chronological evolution of OpenIE technologies and their alignment with significant milestones in NLP development. Consequently, we have not covered multi-domain and multi-lingual datasets or methodologies extensively. While we do address some non-English datasets, specifically Mandarin, and briefly mention multilingual models in Appendix B and model applications across various domains in Appendix C.3, these discussions are not the focal point of our analysis. This limitation is intentional in order to maintain a clear focus on the historical progression of the field rather than the breadth of dataset diversity or the adaptability of methodologies across languages and domains.
Another potential limitation is our survey’s emphasis on the macro aspects of the OpenIE field rather than detailed, micro-level analysis of specific methodologies. As outlined in Section 1, many existing surveys already cover methodologies and models from the pre-LLM era, and we felt that redundant elaboration on these would not add significant value. Post-LLM, despite substantial research leveraging LLMs for traditional IE tasks, there is still a scarcity of studies specifically applying LLMs to OpenIE tasks. This scarcity has constrained our ability to conduct an in-depth survey focused exclusively on LLM methodologies within OpenIE. Nonetheless, from the existing work on LLMs in traditional IE and UIE, detailed in Appendix C, we observe emerging trends that warrant a macro-level analysis. Our approach of integrating and reviewing the field through a historical lens is essential to provide a comprehensive view, enabling a clearer understanding of the task and aiding in the development of a more defined future roadmap.
@@ Acknowledgements
The work was supported by Alibaba Innovative Research (AIR) project support funding. We thank Yafu Li and all reviewers for their generous help and advice during this research.
Appendix A Table of OpenIE Models
OIE16 Re-OIE16 CaRB FewRel TACRED
OpenIE System F1 AUC F1 AUC F1 AUC ARI $B^{3}$ V ARI $B^{3}$ V
Pre-Neural (ORTE) Ollie Schmitz et al. (2012) 38.6 20.2 49.5 31.3 41.1 22.4 - - - - - -
ClausIE Del Corro and Gemulla (2013) 58.0 36.4 64.2 46.4 44.9 22.4 - - - - - -
OpenIE4 Mausam (2016) 58.8 40.8 68.3 50.9 51.6 29.5 - - - - - -
PropS Stanovsky and Dagan (2016) 54.4 32.0 64.2 43.3 31.9 12.6 - - - - - -
Neural Era (ORTE) RnnOIE Stanovsky et al. (2018) 62.0 48.0 - - 49.0 26.1 - - - - - -
OpenIE6 Kolluru et al. (2020a) - - - - 52.7 33.7 - - - - - -
SpanOIE Zhan and Zhao (2020) 69.4 49.1 77.0 65.8 48.5 - - - - - - -
IMoJIE Kolluru et al. (2020b) - - - - 53.5 33.3 - - - - - -
MacroIE Bowen et al. (2021) - - - - 54.8 36.3 - - - - - -
DetIELSOIE Vasilkovsky et al. (2022) - - - - 43.0 27.2 - - - - - -
DetIEIMoJIE Vasilkovsky et al. (2022) - - - - 52.1 36.7 - - - - - -
SMiLe-OIE Dong et al. (2022) - - - - 53.8 34.9 - - - - - -
Neural Era (ORSE) $\textrm{Multi}^{2}\textrm{OIE}$ Ro et al. (2020) - - 83.9 74.6 52.3 32.6 - - - - - -
Gen2OIE Kolluru et al. (2022) - - - - 54.4 32.3 - - - - - -
Gen2OIE (label-rescore) - - - - 54.5 38.9 - - - - - -
OIE@OIA Wang et al. (2022e) 71.6 54.3 85.3 76.9 51.1 33.9 - - - - - -
DragonIE Yu et al. (2022) - - - - 55.1 36.4 - - - - - -
ChunkOIE(SaC-OIA-SP) Dong et al. (2023) - - - - 53.6 35.5 - - - - - -
ChunkOIE(SaC-CoNLL) - - - - 53.2 34.7 - - - - - -
Neural Era (ORC) RSN Wu et al. (2019) - - - - - - 45.3 58.9 70.8 45.9 63.1 64.3
RSN-CV Wu et al. (2019) - - - - - - 54.2 63.8 72.4 - - -
SelfORE Hu et al. (2020) - - - - - - 64.7 67.8 78.3 44.7 54.1 61.9
RSN-BERT Zhao et al. (2021) - - - - - - 53.2 70.9 78.1 75.6 83.4 85.9
RoCORE Zhao et al. (2021) - - - - - - 70.9 79.6 86 81.2 86 88.8
OHRE Zhang et al. (2021b) - - - - - - 64.2 70.5 76.7 - - -
MatchPrompt Wang et al. (2022c) - - - - - - 66.5 72.3 82.2 75.3 83.0 84.5
PromptORE Genest et al. (2022) - - - - - - 43.4 48.8 71.8 - - -
CaPL Duan et al. (2022) - - - - - - 79.4 81.9 88.9 82.9 87.3 89.8
ASCORE Zhao et al. (2023) - - - - - - 67.6 73.5 83.5 78.1 78 83.1
LLM Era IELM GPT-2XL Wang et al. (2022b) - - 35.0 - 22.7 - - - - - - -
GPT-3.5-Turbo ICL Ling et al. (2023) 65.1 - 67.9 - 52.1 - - - - - - -
ChatGPT $n$-shot Qi et al. (2023a) - - - - 55.3 - - - - - - -
Table 4: Performance of OpenIE models. Only report F1 scores for $B^{3}$ and V.
The performances of up-to-date open IE models are summarized in Table 4. The models are categorized the same as that in Section 5 into five groups. For models of pre-neural (ORTE), neural era (ORTE), neural era (ORSE), and LLM era, the F1 scores and accuracy on OIE16, Re-OIE16, and CaRB datasets are reported. For models of neural era (ORC), the ARI, F1 scores for $B^{3}$, and F1 scores for V-measure on FewRel and TACRED datasets are reported.
Appendix B Open IE Methodologies in Details
B.1 Open Relation Triplet Extraction
B.1.1 Labeling
@@
B.2 Open Relation Span Extraction
@@
B.3 Open Relation Clustering
@@
B.4 Neural Model Era: Other Settings
@@
Appendix C LLMs for IE in general
@@
C.1 Universal Information Extraction
@@
Code-Based Schema. Despite their empirical success, natural language-based approaches face challenges in generating outputs for IE tasks due to the distinct syntax and structure that differ from the training data of LLMs Bi et al. (2024). In response to these limitations and leveraging recent advancements in Code-LLMs Chen et al. (2021b), researchers have begun to utilize Code-LLMs for structure generation tasks Wang et al. (2022f), as code, a formalized language, adeptly describes structural knowledge across various schemas universally Guo et al. (2023b). For instance, Li et al. (2023d) present CodeIE, which translates structured prediction tasks such as NER and RE into code generation, employing Python functions to create task-specific schemas and using few-shot learning to instruct Code-LLMs. Guo et al. (2023b) introduce Code4UIE, utilizing Python classes to define task-specific schemas for diverse structural knowledge universally. Similarly, Sainz et al. (2023) propose GoLLIE, which employs Python classes to encode IE tasks and, in addition, integrates task-specific guidelines as docstrings, enhancing the robustness of fine-tuned Code-LLMs to schemas not encountered during training.
C.2 Role of LLMs in IE System
@@
Data Synthesis. A prominent application of LLMs in IE systems is the synthesis of high-quality training data, as data curation through human annotation is time-consuming and labor-intensive. One approach employs LLMs as annotators within a learning loop Zhang et al. (2023a), while another strategy involves using LLMs to inversely generate natural language text from structured data inputs Josifoski et al. (2023); Ma et al. (2023), thereby producing large-scale, high-quality training data for IE tasks.
@@ Appendix D Datasets
Synthesized datasets using LLMs on IE expands significantly compared to previous ones in both the size of the datasets and data qualities. While the methodologies for synthesizing these datasets have been extensively explored within the domain of closed Information Extraction (ClosedIE) Zhang et al. (2023b); Xu et al. (2023a), where researchers claims the proposed methods can be adapted for OpenIE setting Josifoski et al. (2023), there remains a notable gap in the literature regarding comprehensive studies on synthesized datasets for OpenIE.
Appendix E Evaluation
@@
Appendix F Source of Information
@@
F.1 Input-based Information
@@ F.2 External Knowledge
Hierarchical information that implicitly exists in languages, which can be explicitly exhibited by knowledge bases, benefits knowledge representation learning Wang et al. (2014); Lin et al. (2015); Hu et al. (2015); Xie et al. (2016). In addition, KBs contain fine-grained factual knowledge that provides background information and hierarchical structures needed for relation extraction. Compared to traditional clustering, KB can provide hierarchical information that helps represent and cluster relations in a more organized way Zhang et al. (2021a) and hierarchical factual knowledge for data augmentation Fangchao et al. (2021).

### v7

@@ 6.3 Future Directions
Although we see the momentum of blurred gaps between OpenIE and standard IE with the impact of LLM, the fundamental task itself remains. Then how can traditional OpenIE research inspire IE research in the LLM era? Following we discuss future directions draw from reflections on a chronological perspective.
OpenIE datasets are growing but remain small and narrow in scope. Insights from traditional OpenIE research suggest that future expansions are needed to include more languages, domains, and broader sources. LLMs offer the opportunity to improve this through their capabilities in synthesizing and augmenting data. While synthesized datasets have been extensively explored within the domain of standard IE Zhang et al. (2023a); Xu et al. (2023a), with researchers claiming that the proposed methods can be adapted for OpenIE Josifoski et al. (2023), there is a notable gap regarding comprehensive studies on synthesized datasets for OpenIE. Addressing this gap could facilitate the creation of cross-domain datasets and the integration of existing datasets and tasks.
@@
As discussed in Section 6.2, LLMs enable the exploration of various IE tasks with universal frameworks (UIE). Despite advances, most LLM-based UIE systems focus on standard IE tasks and often overlook OpenIE, a complex challenge within the IE spectrum. LLMs are inherently suited for OpenIE due to their extensive pre-trained knowledge. Therefore, the primary challenge of LLMs lies not in extracting relational information but in accurately interpreting and following task-specific instructions, as well as mitigating hallucination. Integrating traditional approaches into LLM-based frameworks might address these current shortcomings of LLMs. Additionally, these approaches can provide insights for developing more robust, faithful, and reliable fundamental LLMs.
More comprehensive automatic metrics are needed to evaluate LLM-based approaches. As discussed in Sections 2 and 4, task settings and corresponding evaluation metrics develop hand-in-hand. Now the changes brought by LLM calls for a more holistic and update-to-date evaluation metrics in many NLP tasks Gong et al. (2024); Ai et al. (2024a, c, b). Current efforts explore different options, as noted in Section 4 and Appendix D, but aspects such as faithfulness still rely heavily on human evaluation and lack a commonly accepted metric. Developing new, comprehensive automatic evaluation methods that capture nuanced aspects of OpenIE output, such as semantic coherence, factual accuracy, and information completeness, will lead to more robust and reliable LLM-based OpenIE systems. These metrics can address the unique challenges posed by LLMs, including their propensity for generating diverse and open-ended outputs.
Multilinguality is critical, yet there is a noticeable gap in multilingual capabilities. Although the models discussed in Appendix A.4 show efforts to adapt across languages, there is a notable deficiency in initiatives using LLMs to tackle these challenges. This highlights the need to develop datasets, metrics, and robust benchmarks, particularly for low-resource languages, to enhance the capabilities of multilingual OpenIE.
Latency, cost, and distillation. Reviewing the development of models for OpenIE, we see the trend that recent development in LLMs introduce a more expensive system with higher latency, especially using close sourced LLMs such as GPT. Although the rapid iteration of models shows cuts on cost and latency, more effective solutions could be possible with knowledge distillation from LLMs onto specialized SLMs, revisiting the prior neural model era we discussed.
7 Conclusion
We systematically survey the development of OpenIE from a chronological perspective, highlighting historical trends in task settings and model development. We draw important connections and derive lessons from the influence of technology on task settings, examining the advantages and disadvantages of both past and present methods. Furthermore, we explore the increasingly blurred distinctions between OpenIE and standard IE. For researchers in LLMs, past work should not be overlooked; instead, it should serve as a valuable resource for future inquiries.
@@ 7 Conclusion
Our survey primarily concentrates on the chronological evolution of OpenIE technologies and their alignment with significant milestones in NLP development. Consequently, we have not covered multi-domain and multi-lingual datasets or methodologies extensively. While we do address some non-English datasets, specifically Mandarin, and briefly mention multilingual models in Appendix A and model applications across various domains in Appendix B.3, these discussions are not the focal point of our analysis. This limitation is intentional in order to maintain a clear focus on the historical progression of the field rather than the breadth of dataset diversity or the adaptability of methodologies across languages and domains.
Another potential limitation is our survey’s emphasis on the macro aspects of the OpenIE field rather than detailed, micro-level analysis of specific methodologies. As outlined in Section 1, many existing surveys already cover methodologies and models from the pre-LLM era, and we felt that redundant elaboration on these would not add significant value. Post-LLM, despite substantial research leveraging LLMs for standard IE tasks, there is still a scarcity of studies specifically applying LLMs to OpenIE tasks. This scarcity has constrained our ability to conduct an in-depth survey focused exclusively on LLM methodologies within OpenIE. Nonetheless, from the existing work on LLMs in standard IE and UIE, detailed in Appendix B, we observe emerging trends that warrant a macro-level analysis. Our approach of integrating and reviewing the field through a historical lens is essential to provide a comprehensive view, enabling a clearer understanding of the task and aiding in the development of a more defined future roadmap.
@@ Acknowledgements
The work was supported in part by Alibaba Innovative Research (AIR) project support funding. We thank Yafu Li and all reviewers for their generous help and advice during this research. This work was supported in part by the US Department of Defense under the DARPA CCU program. Any opinions expressed herein are those of the authors and do not necessarily reflect the views of the U.S. Department of Defense or the U.S. Government.
Appendix A Open IE Methodologies in Details
A Chronological Overview of Open IE methods are summarized in Figure 3.
A.1 Open Relation Triplet Extraction
A.1.1 Labeling
@@
A.2 Open Relation Span Extraction
@@
A.3 Open Relation Clustering
@@
A.4 Neural Model Era: Other Settings
@@
Appendix B LLMs for IE in general
@@
B.1 Universal Information Extraction
@@
Code-Based Schema. Despite their empirical success, natural language-based approaches face challenges in generating outputs for IE tasks due to the distinct syntax and structure that differ from the training data of LLMs Bi et al. (2024). In response to these limitations and leveraging recent advancements in Code-LLMs Chen et al. (2021), researchers have begun to utilize Code-LLMs for structure generation tasks Wang et al. (2022e), as code, a formalized language, adeptly describes structural knowledge across various schemas universally Guo et al. (2023b). For instance, Li et al. (2023d) present CodeIE, which translates structured prediction tasks such as NER and RE into code generation, employing Python functions to create task-specific schemas and using few-shot learning to instruct Code-LLMs. Guo et al. (2023b) introduce Code4UIE, utilizing Python classes to define task-specific schemas for diverse structural knowledge universally. Similarly, Sainz et al. (2023) propose GoLLIE, which employs Python classes to encode IE tasks and, in addition, integrates task-specific guidelines as docstrings, enhancing the robustness of fine-tuned Code-LLMs to schemas not encountered during training.
B.2 Role of LLMs in IE System
@@
Data Synthesis. A prominent application of LLMs in IE systems is the synthesis of high-quality training data, as data curation through human annotation is time-consuming and labor-intensive. One approach employs LLMs as annotators within a learning loop Zhang et al. (2023b), while another strategy involves using LLMs to inversely generate natural language text from structured data inputs Josifoski et al. (2023); Ma et al. (2023), thereby producing large-scale, high-quality training data for IE tasks.
@@ Appendix C Datasets
Synthesized datasets using LLMs on IE expands significantly compared to previous ones in both the size of the datasets and data qualities. While the methodologies for synthesizing these datasets have been extensively explored within the domain of closed Information Extraction (ClosedIE) Zhang et al. (2023a); Xu et al. (2023a), where researchers claims the proposed methods can be adapted for OpenIE setting Josifoski et al. (2023), there remains a notable gap in the literature regarding comprehensive studies on synthesized datasets for OpenIE.
Appendix D Evaluation
@@
D.1 Evaluation Metrics of ORC
$B^{3}$ Metric
$B^{3}$ is an instance-based method that computes the precision and recall for each instance in the test set by comparing the cluster containing the instance in the prediction results with the cluster containing the instance in the ground truth (golden answer). $B^{3}$ averages the precision and recall for each instance and then computes the harmonic mean to provide a final score.
Precision for an element $i$:
$P(i)=\frac{|C_{i}\cap T_{i}|}{|C_{i}|}$
Where: $C_{i}$ is the set of elements in the same cluster as $i$. $T_{i}$ is the set of elements in the same true class as $i$.
Recall for an element $i$:
$R(i)=\frac{|C_{i}\cap T_{i}|}{|T_{i}|}$
Average Precision and Recall:
$P=\frac{1}{N}\sum_{i=1}^{N}P(i),\quad R=\frac{1}{N}\sum_{i=1}^{N}R(i)$
Where $N$ is the total number of elements.
F1 Score:
$F1=2\cdot\frac{P\cdot R}{P+R}$
V-measure
V-measure is an entropy-based, instance-based clustering evaluation metric that introduces conditional entropy, requiring higher purity of clusters. Compared to $B^{3}$, the existence of a few incorrect instances in a relatively pure cluster decreases the score more significantly, thereby punishing clustering results more harshly. The V-measure F1 score is the harmonic mean of homogeneity and completeness.
Homogeneity $h$:
$h=1-\frac{H(C\mid K)}{H(C)}$
Where $H(C\mid K)$ is the conditional entropy of the classes given the cluster assignments; $H(C)$ is the entropy of the classes.
Completeness $c$:
$c=1-\frac{H(K\mid C)}{H(K)}$
Where $H(K\mid C)$ is the conditional entropy of the clusters given the class assignments; $H(K)$ is the entropy of the clusters.
V-measure:
$V=\frac{2\cdot h\cdot c}{h+c}$
Adjusted Rand Index (ARI)
Adjusted Rand Index (ARI) measures the similarity between predicted and true clusterings by counting all pair-wise assignments in the same or different groups. It adjusts for chance groupings, with a threshold ranging from 0 to 1, representing random groupings to perfectly accurate groupings. Compared to $B^{3}$ and V-measure, ARI is less sensitive to extreme values like precision or homogeneity, providing a more balanced evaluation.
Rand Index (RI):
$RI=\frac{a+b}{a+b+c+d}$
Where $a$ is the number of pairs of elements that are in the same set in both the predicted and true clusterings, $b$ is the number of pairs of elements in different sets in both the predicted and true clusterings, $c$ is the number of pairs of elements in the same set in the true clustering but not in the predicted clustering, and $d$ is the number of pairs of elements in the same set in the predicted clustering but not in the true clustering.
Adjusted Rand Index (ARI):
$ARI=\frac{\sum_{ij}\binom{n_{ij}}{2}-\left[\sum_{i}\binom{a_{i}}{2}\sum_{j}\binom{b_{j}}{2}\right]/\binom{n}{2}}{\frac{1}{2}\left[\sum_{i}\binom{a_{i}}{2}+\sum_{j}\binom{b_{j}}{2}\right]-\left[\sum_{i}\binom{a_{i}}{2}\sum_{j}\binom{b_{j}}{2}\right]/\binom{n}{2}}$
Where $n_{ij}$ is the number of elements in both predicted cluster $i$ and true cluster $j$.
Appendix E Source of Information
@@
E.1 Input-based Information
@@ E.2 External Knowledge
Hierarchical information that implicitly exists in languages, which can be explicitly exhibited by knowledge bases, benefits knowledge representation learning Wang et al. (2014); Lin et al. (2015); Hu et al. (2015); Xie et al. (2016). In addition, KBs contain fine-grained factual knowledge that provides background information and hierarchical structures needed for relation extraction. Compared to traditional clustering, KB can provide hierarchical information that helps represent and cluster relations in a more organized way Zhang et al. (2021b) and hierarchical factual knowledge for data augmentation Fangchao et al. (2021).
@@
Appendix F Table of Traditional OpenIE Models
Model Method Source of Information
Task
Setting
Dataset Evaluation (Result)
TextRunner
Banko et al. (2007)
Dependency Parser, NP Chunker,
CRF, Naive Bayes Classifier
syntactic, dependency ORTE 400 Web Average Error Rate (12%)
WOE
Wu and Weld (2010)
TextRunner,
Self-supervised Learning
syntactic, dependency ORTE
300 news
300 Wikipedia
300 Web
Precision-Recall Curve
ReVerb
Fader et al. (2011)
Syntactic Constraints,
Lexical Contraints, CRF
syntactic, dependency ORTE 500 Web
Precision-Recall Curve,
AUC (1.3*$\textrm{WOE}^{parse}$, 2*TextRunner)
Ollie
Schmitz et al. (2012)
ReVerb, Bootstrap,
Open Pattern Learning
syntactic, dependency ORTE
300 news (from WOE)
300 Wikipedia (from WOE)
300 biology
Precision-Yield Curve,
AUC (1.9*$\textrm{WOE}^{parse}$, 2.7*ReVerb)
OpenIE4
Mausam (2016)
SrlIE Christensen et al. (2011),
RelNoun Pal et al. (2016)
syntactic, dependency ORTE Not Reported
Precision-Yield Curve,
AUC (1.32*Ollie, 4*ReVerb)
ClausIE
Del Corro and Gemulla (2013)
Dependency Parser,
Clause-based Model
syntactic, dependency ORTE
500 Web (from REVERB)
200 Wikipedia
200 news
Precision-Yield Curve,
# of correct extractions / # of extractions
RnnOIE Stanovsky et al. (2018) Bi-LSTM, Softmax word emb, POS emb ORTE OIE2016 AUC (48), F1 (62)
WEB AUC (47), F1 (67)
NYT AUC (25), F1 (35)
PENN AUC (26), F1 (44)
NeuralOIE
Cui et al. (2018)
LSTM, Copy Attention word emb ORTE OIE2016 AUC (27)
IMoJIE
Kolluru et al. (2020b)
BERT, LSTM, CopyAttention word emb ORTE CaRB AUC (33.3), F1 (53.5)
SpanOIE Zhan and Zhao (2020) Bi-LSTM, Span-consistent Greedy Search word emb, POS emb, dependency relation emb ORTE OIE2016 AUC (48.9), F1 (68.65)
Re-OIE2016 AUC (65.9), F1 (78.50)
$\textrm{Multi}^{2}\textrm{OIE}$ Ro et al. (2020) BERT, Multihead Attention word emb, position emb, avg vector of predicates ORTE Re-OIE2016 AUC (74.6), F1 (83.9)
CaRB AUC (32.6), F1 (52.3)
OpenIE6
Kolluru et al. (2020a)
Iterative Grid Labeling,
BERT, Self-attention
word emb,
dependency feature
ORTE CaRB AUC (33.7), F1 (52.7)
HNN4ORT Jia et al. (2022) ON-LSTM, CNN, Attention word emb, POS emb, argument emb, local/global features ORTE Wikipedia F1 (79.8)
NYT F1 (74.5)
Reverb F1 (81.7)
UORE
Elsahar et al. (2017)
Re-weight Word Emb,
TF-IDF, PCA, HAC
word emb,
dependency
ORC NYT-FB F1 (41.6)
RSN
Wu et al. (2019)
Relational Siamese Network,
CNN, HAC, Louvain
word emb ORC FewRel ${B}^{3}$: P (48.9) R (77.5) F1 (59.9)
SelfORE Hu et al. (2020) Bootstrapping Self-supervision, BERT, K-means, Adaptive Clustering word emb ORC NYT+FB
ARI (40.3),
${B}^{3}$: P (49.1) R (47.3) F1 (51.1),
V: F1 (46.6) Hom (45.7) Comp (47.6)
T-REx SPO
ARI (33.7),
${B}^{3}$: P (41.0) R (39.4) F1 (42.8),
V: F1 (41.4) Hom (40.3) Comp (42.5)
T-REx DS
ARI (20.1),
${B}^{3}$: P (32.9) R (29.7) F1 (36.8),
V: F1 (32.4) Hom (30.1) Comp (35.1)
OHRE Zhang et al. (2021b) CNN, Virtual Adversarial Training, Reconstruction Loss, Dynamic Hierarchical Triplet Loss, Louvain word emb, hierarchical information ORC FewRel Hierarchy
ARI (64.2),
${B}^{3}$: P (64.5) R (77.7) F1 (70.5),
V: F1 (76.7) Hom (73.8) Comp (79.9)
NYT-FB Hierarchy
ARI (31.9),
${B}^{3}$: P (31.4) R (72.3) F1 (43.8),
V: F1 (60.0) Hom (49.9) Comp (75.3)
ElementORE Fangchao et al. (2021) BERT, T5 Raffel et al. (2020), Structure Causal Model, K-means word emb, hierarchical information ORC T-REx SPO
ARI (36.6 ),
${B}^{3}$: P (46.7) R (43.4) F1 (45.0),
V: F1 (45.3) Hom (45.4) Comp (45.2)
T-REx DS
ARI (25.0),
${B}^{3}$: P (40.2) R (45.9) F1 (42.9),
V: F1 (47.3) Hom (46.9) Comp (47.8)
RoCORE
Zhao et al. (2021)
Relation-oriented Representation,
BERT, K-means
word emb ORC FewRel
ARI (70.9),
${B}^{3}$: P (75.2) R (84.6) F1 (79.6),
V: F1 (86.0) Hom (83.8) Comp (88.3)
DeepEx Wang et al. (2021) BERT, Attention, Beam Search, Contrastive Pre-training NP chunks, word emb, triplet emb ORSE OIE2016 AUC (58.6), F1 (72.6)
WEB AUC (82.4), F1 (91.2)
NYT AUC (72.5), F1 (85.5)
PENN AUC (81.5), F1 (88.5)
Table 4: Milestone and representative models of pre-LLM era. ("V" denotes "V-measure", and "emb" stands for "embedding".)